OpenStudy (anonymous):

Solve sin2x - 2cos^2x=0

OpenStudy (anonymous):

\[\sin 2x = 2\sin(x)\cos(x)\]\[2\sin(x)\cos(x) - 2\cos^2(x) = 0\]\[2\cos(x)(\sin(x)-\cos(x)) = 0\]\[\cos(x) = 0, x = 90, 180\]\[\tan(x) = 1, x = 45, 225\]

OpenStudy (anonymous):

So x = 45, 90, 180, 225

OpenStudy (anonymous):

where did the tan(x) come from?

OpenStudy (anonymous):

45, 90, and 225 are correct but the answer doesn't include 180 , but include 3pi/4 which is 135?